OpenStudy (anonymous):

Give the values of a, b, and c from the general form of the equation (2x + 1)(x - 2) = 0. a) a = 2, b = -3, c = -2 b) a = 2, b = 5, c = -2 c) a = 3, b = 1, c = -1

OpenStudy (anonymous):

i guess they want you to multiply \[(2x+1)(x-2)\] do you know how to do that?

OpenStudy (anonymous):

you have to do four multiplications, which some idiot math teachers like to call "foil" but in fact is the distributive law \[(2x+1)(x-2)=2x\times x-2\times 2x+1\times x+1\times (-2)\] which cleans up to \[2x^2-4x+x-2\] combining like terms gives \[x^2-3x-1\]

OpenStudy (anonymous):

another typo it must be my bed time last line should be \[\huge 2x^2-3x-2\]